# Uncertainty Quantification in Retrieval Augmented Question Answering

### v1

@@ Abstract
Uncertainty Quantification in Retrieval Augmented Question Answering
Laura Perez-Beltrachini  and Mirella Lapata Institute for Language, Cognition and Computation School of Informatics, University of Edinburgh 10 Crichton Street, Edinburgh EH8 9AB {lperez,mlap}@inf.ed.ac.uk
@@ 1 Introduction
Retrieval augmented Question Answering (QA) allows QA models to overcome knowledge gaps at test time through access to evidence in the form of retrieved passages (Lewis et al., 2020; Guu et al., 2020; Izacard et al., 2024). Recent work leverages external retrievers (Chen et al., 2017; Izacard and Grave, 2021b) and the language understanding and generation capabilities of Large Language Models (LLMs; Brown et al. 2020; Ouyang et al. 2024) to predict answers based on questions and retrieved passages which are provided as input context. In Figure 1, we show an example of a question (Who sings Does He Love Me with Reba?), retrieved passages, and predicted answers.
@@
Linda Davis Does He Love You. Does He Love You "Does He Love You" is a song written by Sandy Knox and Billy Stritch, and recorded as a duet by American country music artists Reba McEntire and Linda Davis. It was released in August 1993 as the first single from Reba’s album "Greatest Hits Volume Two". It is one of country music’s several songs [cont.] 4.1
Reba McIntyre and Brooks & Dunn’s Ronnie Dunn Reba: Duets. The first collaborator on the album was LeAnn Rimes, who recorded the track, "When You Love Someone Like That" which also -3.94 appeared on LeAnn Rimes’s "Family" album that same year. Jurek called the duet between the pair "stellar," while "about.com" called the pairing "an undeniable outcome of perfection. Reba’s strong country voice with LeAnn’s young, soulful sound [cont.] -3.94
Trisha Yearwood Reba: Duets. Artist, Trisha Yearwood on the song, "She Can’t Save Him", which was formerly released as a single by Canadian country artist, Lisa -3.91 Brokop. Tracks six and seven were collaborations with American pop artist, Carole King and country artist, Kenny Chesney, who both help in providing musical variations towards [cont.] -3.91
@@
Figure 1: Example question from Natural Questions dataset Kwiatkowski et al. (2019) with three top-retrieved passages using Contriever-MSMARCO (Izacard et al., 2022). On top of each passage, we show the answer generated by Gemma2-9B when prompted with that passage and the question. The QA model answers correctly (green) only when prompted with the first passage. Numbers at the bottom right of each passage are utility scores predicted by our model (higher values indicate more useful passages).
Retrieval augmented QA architectures have proven beneficial in increasing LLM performance on tail knowledge (Izacard et al., 2024; Mallen et al., 2023), reducing hallucinations in the generated answers, and even improving model calibration (Jiang et al., 2021). However, there are various ways in which retrieval augmented QA can go wrong at inference time. The set of retrieved passages is far from perfect (Sciavolino et al., 2021; Yoran et al., 2024; Kasai et al., 2024) containing irrelevant or misleading evidence, the model might be under-trained to read certain passages and reason over these and the question (Izacard et al., 2024; Liu et al., 2024b), and the question can be ambiguous or unanswerable (Kasai et al., 2024). In such cases of uncertainty, QA models should ideally be able to deal with it (e.g., communicating it or abstaining from answering) rather than risking an incorrect response.
A good deal of previous work has focused on quantifying answer uncertainty in the context of closed-book QA tasks, where the answer is predicted based on a question and the model’s encoded knowledge. Sampling-based methods rely on output discrepancies among multiple predictors of the same input (Gal and Ghahramani, 2016; Lakshminarayanan et al., 2017). They measure diversity on a set of answers (Kuhn et al., 2023; Chen and Mueller, 2024) sampled via temperature scaling (Guo et al., 2017), with larger variance indicating higher uncertainty. LLM-based methods rely on the QA model’s own judgment of uncertainty (Kadavath et al., 2022; Lin et al., 2022; Tian et al., 2023). Through prompting, the model is encouraged to express its uncertainty (e.g., 0.5 or ‘almost certain’), either alongside the predicted answer (Lin et al., 2022; Tian et al., 2023) or after generating it Kadavath et al. (2022); Tian et al. (2023).
@@
In this paper, we focus on retrieval augmented QA and hypothesize that a passage is useful, if a model can correctly answer questions based on it. If passages are informative and prime the QA model towards appropriate knowledge, we expect it to produce a correct answer. On the contrary, if passages are misleading and the question falls outside the QA model’s knowledge, it is likely to produce an incorrect answer — either factually inaccurate or entirely fabricated. We quantify the utility of a retrieved passage with a small neural model trained on utility judgements predicted by the target QA model. We borrow ideas from direct uncertainty quantification approaches (Van Amersfoort et al., 2020; Lahlou et al., 2023) but do not decompose uncertainty or outline shifts in the input distribution. We make utility predictions for each retrieved passage which we then use to estimate the uncertainty of the QA model.
@@
We evaluate our approach on short-form question answering tasks (see Figure 1 for an example). Results on six datasets show that our uncertainty estimator outperforms existing sampling-based methods (especially in complex reasoning questions and adversarial QA settings with rare entities or unanswerable questions) while being more test-time efficient. Sampling-based solutions are expensive for in-production QA systems, in terms of latency (see comparison in Appendix A) and cost (e.g., QA engines built on top of proprietary language models would need to process relatively long prompts). Our contributions can be summarized as follows:
@@
  • We propose to quantify QA model uncertainty via estimating the utility of the passages it is provided with.
@@
  • We (contrastively) train a small neural model on utility scores obtained through combining accuracy (is the generated answer correct?) and entailment (is the answer supported by the passage?) metrics.
@@
  • We show that utility scores predicted by our model can further improve QA accuracy by re-ranking passages obtained via an external retriever (Liu et al., 2024b).
@@ 2 Related Work
Several methods have been proposed to predict answer uncertainty in QA; however, none of them has analysed uncertainty in retrieval augmented QA models. Many existing approaches rely on the assumption that output variation is an expression of model uncertainty (Kuhn et al., 2023; Farquhar et al., 2024; Chen and Mueller, 2024). For example, Kuhn et al. (2023) first cluster answers with similar meaning (in a sample) via natural language inference before computing entropy while Chen and Mueller (2024) focus on black-box models; they also compute similarities in the set of answers but associate them with a model self-judgement of confidence. These approaches are expensive to run at inference time for a production QA system, they require several inference steps in addition to performing similarity computations which can become more complex with longer answers (Zhang et al., 2024). Hou et al. (2024) focus on quantifying aleatoric uncertainty (i.e., uncertainty in the data) caused by ambiguous questions, an approach which could be combined with ours.
Judging the Utility of Retrieved Passages
@@
Previous work has analysed the quality of retrieved passages (Yu et al., 2023; Asai et al., 2024; Wang et al., 2024; Xu et al., 2024; Yoran et al., 2024) as they can be irrelevant or misleading. Asai et al. (2024) make use of an external critic model to judge whether a question requires retrieval (or not) and whether the retrieved passages are relevant to formulate the answer. While they analyse passage relevance, this decision is taken by an external extreme-scale critic (e.g., GPT-4) and used to fine-tune their QA model. In contrast, we elicit utility judgements from the target QA model and use these to train a secondary small-scale model to predict passage utility (i.e., our approach does not require LLM fine-tuning). Other work creates auxiliary tasks around retrieved passages enforcing the QA model to reason on them; e.g., by taking notes about each passage (Yu et al., 2023) or generating passage summaries (Xu et al., 2024). These methods also use extreme-scale LLMs to generate training data for fine-tuning a retrieval augmented QA model. Park et al. (2024) select in-context examples with conflicting content (e.g., different dates for a given event) in order to improve LLM reasoning on input passages. These approaches aim at improving QA performance while our primary goal is modelling QA uncertainty.
Improving Retrieval via QA Performance
Previous work has focused on jointly training the retriever and QA modules end-to-end (Lee et al., 2019; Lewis et al., 2020; Izacard and Grave, 2021a). This joint training scheme is very expensive for current (extreme-scale) LLMs. Our approach can be seen as an intermediate module between the QA model and the external retriever and could be used to provide feedback (i.e., utility scores) for fine-tuning the retriever, however, we leave this to future work. Salemi and Zamani (2024) evaluate the quality of retrieval on QA tasks and show that external judgements (e.g., query-document relevance labels) of passage utility correlate poorly with QA performance.
@@
Some approaches train a specific model to predict answer confidence scores (Dong et al., 2018; Kamath et al., 2020; Zhang et al., 2021; Mielke et al., 2022) by incorporating various features from the input and model output. Our approach predicts answer uncertainty directly from individual passage utilities but could be combined with other features (e.g., output sequence probability). Some work Kamath et al. (2020); Zhang et al. (2021) predicts answer correctness in the context of Reading Comprehension (the task of generating an answer based on a single supportive passage). However, as there is no retrieval involved, the input passage is by default useful, and the main goal is to detect answer uncertainty due to the QA model being under-trained. In our setting, the number and utility of passages varies leading to additional sources of uncertainty (e.g., misleading information).
Our passage utility predictor is related to methods aiming to estimate error directly Lahlou et al. (2023), e.g., by training a secondary model to estimate target model loss; instead, our predictor is trained with sequence-level metrics, i.e., accuracy and entailment, which measure error indirectly.
3 Modelling Answer Uncertainty
@@
When a retrieved passage is useful to answer a given question (such as the first passage in Figure 1 for the question Who sings Does He Love Me with Reba?), the QA model is likely to be confident when generating the answer (Linda Davis). When the passage is not useful (such as the third passage in Figure 1), the QA model is likely to be uncertain and provide an incorrect answer (Trisha Yearwood). Our hypothesis is that the utility of each passage $p$ in $R$ is indicative of the QA model’s uncertainty in generating $y$, when prompted with $R$. If there are passages in $R$ with high utility (e.g., in Figure 1, the first passage is useful to answer the question), it is likely that the QA model will be confident when generating answer $y$. In contrast, if all passages in $R$ have low utility, it is likely that the QA model will be uncertain when generating the answer.
@@ 3.1 Passage Utility Prediction
Intuitively, a retrieved passage $p$ is useful for a QA model $\mathcal{M}$ if $\mathcal{M}$ can correctly answer question $x$ when prompted with $p$. However, the model’s dependence on $p$ may vary. In some cases, $\mathcal{M}$ may generate the correct answer even if $p$ does not explicitly contain it, instead it positively primes the model to draw upon its memorised knowledge. In Figure 1, the first passage has high utility because the QA model generates a correct answer when prompted with it, and explicitly states that “Linda Davis sings alongside Reba McEntire”. In contrast, the second and third passages, while related to the question’s topic, are not useful. The QA model struggles to answer correctly when prompted with them, suggesting uncertainty. Since these passages do not provide helpful information and lead to incorrect answers, their utility is low.
We estimate the utility of passage $p$ in answering question $x$ for QA model $\mathcal{M}$ by combining two key measures: accuracy and entailment. Accuracy, denoted as $a(y)$, indicates whether the generated answer $y$ is correct, while entailment, denoted as $e(y)$, measures the degree to which $p$ supports $y$. Accuracy is determined by an evaluator $A$, and entailment is assessed using a Natural Language Inference (NLI) classifier model $E$. We define the combined passage utility as $\upsilon_{\mathcal{M}}=(a(y)+e(y))/2$ which ranges between 0 and 1, given that $a$ takes values in $\{0,1\}$ and $e$ falls within the $[0,1]$ interval.
@@
We train the passage utility predictor with a Siamese neural network consisting of a BERT-based encoder (Devlin et al., 2019) followed by a pooling and two MLP layers stacked on top of BERT outputs (Fang et al., 2024). The output layer computes the utility score as $\upsilon_{i}=W_{o}h^{L}+b_{o}$ where $h^{L}$ is the vector representation for $(x,p_{i})$ from the last hidden layer (the $L$-th layer) of the network. At inference time, we compute a single utility score for each passage. We provide implementation and training details in Section 4.
To strengthen the role of accuracy prediction as a training signal and regularise the range of utility values learned by the ranking scheme, we combine the ranking objective in Equation (2) with the following Binary Cross Entropy (BCE; Sculley 2010) objective:
$\mathcal{L}_{BCE}=\sum_{(x,p)\in\{(x,p_{i}),(x,p_{j})\}}[a\times\log(p(x,p))+$ (3)
$(1-a)\times\log(1-p(x,p))],$
@@
Both ranking and BCE objectives are compatible with gold annotations that could be provided in active learning or interactive settings (Simpson et al., 2020; Fang et al., 2024). For example, moderators of the QA system would provide judgments on the accuracy of the answers it predicts (e.g., correct/incorrect) and whether these are supported by the retrieved passages (e.g., best or worse).

### v2

@@ 1 Introduction
Retrieval augmented Question Answering (QA) allows QA models to overcome knowledge gaps at test time through access to evidence in the form of retrieved passages (Lewis et al., 2020; Guu et al., 2020; Izacard et al., 2024). Recent work leverages external retrievers (Chen et al., 2017; Izacard & Grave, 2021) and the language understanding and generation capabilities of Large Language Models (LLMs; Brown et al. 2020; Ouyang et al. 2024) to predict answers based on questions and retrieved passages which are provided as input context. In Figure 1, we show an example of a question (Who sings Does He Love Me with Reba?), retrieved passages, and predicted answers.
Retrieval augmented QA architectures have proven beneficial in increasing LLM performance on tail knowledge (Izacard et al., 2024; Mallen et al., 2023), reducing hallucinations in the generated answers (Shuster et al., 2021), and even improving model calibration (Jiang et al., 2021). However, there are various ways in which retrieval augmented QA can go wrong at inference time. The set of retrieved passages is far from perfect (Sciavolino et al., 2021; Yoran et al., 2024; Kasai et al., 2024) containing irrelevant or misleading evidence, the model might be under-trained to read certain passages and reason over these and the question (Izacard et al., 2024; Liu et al., 2024b), and the question can be ambiguous or unanswerable (Kasai et al., 2024). Ultimately, QA models may not follow the provided passages (Xie et al., 2024; Joren et al., 2025). When faced with uncertainty, QA models should ideally acknowledge it (e.g., by communicating it) rather than risk an incorrect response.
A good deal of previous work has focused on quantifying answer uncertainty in the context of closed-book QA tasks, where the answer is predicted based on a question and the model’s encoded knowledge. Sampling-based methods rely on output discrepancies among multiple predictors of the same input (Gal & Ghahramani, 2016; Lakshminarayanan et al., 2017). They measure diversity on a set of answers (Kuhn et al., 2023; Chen & Mueller, 2024) sampled via temperature scaling (Guo et al., 2017), with larger variance indicating higher uncertainty. LLM-based methods rely on the QA model’s own judgment of uncertainty (Kadavath et al., 2022; Lin et al., 2022; Tian et al., 2023). Through prompting, the model is encouraged to express its uncertainty (e.g., 0.5 or ‘almost certain’), either alongside the predicted answer (Lin et al., 2022; Tian et al., 2023) or after generating it (Kadavath et al., 2022; Tian et al., 2023). None of these approaches has been applied in the context of retrieval augmented QA.
@@
Linda Davis Does He Love You. Does He Love You "Does He Love You" is a song written by Sandy Knox and Billy Stritch, and recorded as a duet by American country music artists Reba McEntire and Linda Davis. It was released in August 1993 as the first single from Reba’s album "Greatest Hits Volume Two". It is one of country music’s several songs [cont.] 2.77
Patti LaBelle Does He Love You. on Patti LaBelle’s album, F̈lame.̈ The song features a vocal battle between two female narrators who are in love with the same man. Both women know that the man is being unfaithful to them and are wondering who he truly loves. The big-budget, Jon Small-directed video was filmed over 3 days in mid-1993. It begins with Reba in her dressing room wearing a lilac feather gown, where she sees a picture of her lover, which she glances at. [cont.] -0.10
@@
Figure 1: Example question from Natural Questions dataset (Kwiatkowski et al., 2019) with two top-retrieved passages using Contriever-MSMARCO (Izacard et al., 2022). On top of each passage, we show the answer generated by Gemma2-9B when prompted with that passage and the question. The QA model answers correctly (green) only when prompted with the first passage. Numbers at the bottom right of each passage are utility scores predicted by our model (higher values indicate more useful passages).
@@
In this paper, we focus on answer uncertainty estimation in the context of retrieval augmented QA. We hypothesize that a passage is useful, if a model can correctly answer questions based on it. If passages are informative and prime the QA model towards appropriate knowledge Geva et al. (2021), we expect it to produce a correct answer. On the contrary, if passages are irrelevant or misleading and the question falls outside the QA model’s knowledge, it is likely to produce an incorrect answer, either factually inaccurate or entirely fabricated. Importantly, this notion of utility is based on how the target QA model will answer with the provided passages and not on what an external judge (e.g., entailment model) thinks about them. We quantify the utility of a retrieved passage with a small neural model trained on utility judgments obtained by observing the target QA model’s answering behavior. We borrow ideas from direct uncertainty quantification approaches (Van Amersfoort et al., 2020; Lahlou et al., 2023) but do not decompose uncertainty or outline shifts in the input distribution. We make utility predictions for each retrieved passage which we then use to estimate the uncertainty of the QA model when prompted with a set thereof.
@@
We evaluate our approach on short-form information-seeking QA tasks (Rodriguez & Boyd-Graber, 2021) (see Figure 1 for an example). Results on six datasets show that our uncertainty estimator is comparable or outperforms existing sampling-based methods while being more test-time efficient. Sampling-based solutions are expensive for in-production QA systems, in terms of latency and cost (e.g., QA engines built on top of proprietary language models would need to process relatively long prompts). Moreover, our experiments reveal that variation is less prominent in model answers in the context of retrieval augmented QA (e.g., the QA model is more confident on incorrect answers supported by retrieved passages in the prompt). Our contributions can be summarized as follows:
@@
  • We quantify QA model uncertainty via estimating the utility of the passages it is provided with.
@@
  • We (contrastively) train a small neural model on utility scores obtained through combining accuracy (is the generated answer correct?) and entailment (is the generated answer supported by the passage?) metrics.
@@ 2 Related Work
Several methods have been proposed to predict answer uncertainty in QA; however, none of them has analysed uncertainty in retrieval augmented QA models. Many existing approaches rely on the assumption that output variation is an expression of model uncertainty (Kuhn et al., 2023; Farquhar et al., 2024; Chen & Mueller, 2024). For example, Kuhn et al. (2023) first cluster answers with similar meaning (in a sample) via natural language inference before computing entropy while Chen & Mueller (2024) focus on black-box models; they also compute similarities in the set of answers but associate them with a model self-judgement of confidence. These approaches are expensive to run at inference time for a production QA system, they require several inference steps in addition to performing similarity computations which can become more complex with longer answers (Zhang et al., 2024b). Hou et al. (2024) focus on quantifying aleatoric uncertainty (i.e., uncertainty in the data) caused by ambiguous questions, an approach which could be combined with ours.
Judging the Quality of Retrieved Passages
@@
Previous work has analysed the quality of retrieved passages (Yu et al., 2023; Asai et al., 2024; Wang et al., 2024; Xu et al., 2024; Yoran et al., 2024) as they can be irrelevant or misleading. Asai et al. (2024) make use of an external critic model to create training data exemplifying cases where a question requires retrieval (or not) and, in the case that retrieval is needed, whether retrieved passages contain the information (or not) to formulate the answer. Note that the usefulness judgment is made by an external critic. Then, a QA model (i.e., the Self-RAG model) is trained on this data to learn to reflect by itself whether passages are supportive and relevant and to predict special tokens indicating this. While it is possible to derive uncertainty from those special tokens’ probabilities, they only reflect Self-RAG’s uncertainty state. Our proposal is more general and aims at predicting answer uncertainty in zero-shot QA models (e.g., instruction fine-tuned LLMs). Thus, using Self-RAG special tokens would be like using an off-the-shelf sophisticated classifier (i.e., a specialized textual entailment) to predict the uncertainty of a zero-shot QA model (Yoran et al., 2024). In contrast, we predict uncertainty by observing the errors of the target QA model (in a zero-shot setting). Other work creates auxiliary tasks around retrieved passages enforcing the QA model to reason on them; e.g., by taking notes about each passage (Yu et al., 2023) or generating passage summaries (Xu et al., 2024). These methods also use extreme-scale LLMs to generate training data for fine-tuning a retrieval augmented QA model. Park et al. (2024) select low and high quality in-context examples in order to instruct the LLM to reason on input passages. Concurrent work by Joren et al. (2025) defines the concept of context sufficiency, i.e., whether the content in the set of retrieved passages is sufficient to answer the question and uses an external judge to assess this. However, the external judgment might not agree with the answering behavior of the target QA model (i.e., the judge may indicate that the context is sufficient but the model may still incorrectly answer). All these approaches aim at improving QA performance while our primary goal is modeling QA uncertainty.
@@
Our passage utility predictor is related to methods aiming to estimate error directly (Lahlou et al., 2023), e.g., by training a secondary model to estimate target model loss; instead, our predictor is trained with sequence-level metrics, i.e., accuracy and entailment, which measure error indirectly. Some work (Kamath et al., 2020; Zhang et al., 2021) predicts answer correctness in the context of Reading Comprehension (the task of generating an answer based on a single supportive passage). However, as there is no retrieval involved, the input passage is by default useful, and the main goal is to detect answer uncertainty due to the QA model being under-trained. In our setting, the number and quality of passages varies leading to additional sources of uncertainty (e.g., misleading information). Some approaches train a specific model to predict answer confidence scores (Dong et al., 2018; Kamath et al., 2020; Zhang et al., 2021; Mielke et al., 2022) by incorporating various features from the input and model output. Our approach predicts answer uncertainty directly from individual passage utilities and its predictions could also be combined with other features (e.g., output sequence probability).
3 Modeling Answer Uncertainty
@@
When a retrieved passage is useful to answer a given question (such as the first passage in Figure 1 for the question Who sings Does He Love Me with Reba?), the QA model is likely to be confident when generating the answer (Linda Davis). When the passage is not useful (such as the second passage in Figure 1), the QA model is likely to be uncertain and provide an incorrect answer (Patti LaBelle). Our hypothesis is that the utility of each passage $p$ in $R$ is indicative of the QA model’s uncertainty in generating $y$, when prompted with $R$. If there are passages in $R$ with high utility (e.g., in Figure 1, the first passage is useful to answer the question), it is likely that the QA model will be confident when generating answer $y$. In contrast, if all passages in $R$ have low utility, it is likely that the QA model will be uncertain when generating the answer.
@@ 3.1 Passage Utility Prediction
Intuitively, a retrieved passage $p$ is useful for a QA model $\mathcal{M}$ if $\mathcal{M}$ can correctly answer question $x$ when prompted with $p$. However, the model’s dependence on $p$ may vary. In some cases, $\mathcal{M}$ may generate the correct answer even if $p$ does not explicitly contain it, instead it positively primes the model to draw upon its memorized knowledge. In Figure 1, the first passage has high utility because the QA model generates a correct answer when prompted with it, and explicitly states that “Linda Davis sings alongside Reba McEntire”. In contrast, the second passage, while related to the question’s topic, is not useful. The QA model struggles to answer correctly when prompted with it, suggesting uncertainty. Since this passage does not provide helpful information and leads to incorrect answers, its utility is low.
We estimate the utility of passage $p$ in answering question $x$ for QA model $\mathcal{M}$ by combining two key measures, accuracy and entailment:
$\upsilon_{\mathcal{M}}=\frac{(a(y)+e(y))}{2}$ (2)
Accuracy, denoted as $a(y)$, indicates whether the generated answer $y$ is correct, while entailment, denoted as $e(y)$, measures the degree to which $p$ supports $y$. Accuracy is determined by an evaluator $A$, and entailment is assessed using a Natural Language Inference (NLI) classifier model $E$. The combined passage utility $\upsilon_{\mathcal{M}}$ ranges between 0 and 1, given that $a$ takes values in $\{0,1\}$ and $e$ falls within the $[0,1]$ interval.
@@
We train the passage utility predictor with a Siamese neural network consisting of a BERT-based encoder (Devlin et al., 2019) followed by pooling and two MLP layers stacked on top of BERT outputs (Fang et al., 2024). The output layer computes the utility score as $\upsilon_{i}=W_{o}h^{L}+b_{o}$ where $h^{L}$ is the vector representation for $(x,p_{i})$ from the last hidden layer (the $L$-th layer) of the network. At inference time, we compute a single utility score for each passage. We provide implementation and training details in Section 4.
To strengthen the role of accuracy prediction as a training signal and regularize the range of utility values learned by the ranking scheme, we combine the ranking objective in Equation (3) with the following Binary Cross Entropy (BCE; Sculley 2010) objective:
$\mathcal{L}_{BCE}=\sum_{(x,p)\in\{(x,p_{i}),(x,p_{j})\}}[a\times\log(p(x,p))+(1-a)\times\log(1-p(x,p))],$ (4)
@@
Both ranking and BCE objectives are compatible with gold annotations that could be provided in active learning or interactive settings (Simpson et al., 2020; Fang et al., 2024). For example, moderators of the QA system would provide judgments on the accuracy of the answers it predicts (e.g., correct/incorrect) and the extent to which these are supported by the retrieved passages (e.g., not supported to fully supported).